find the slope of the graph of the linear function "f". f(-3)=-9,f(3)=9
Found 2 solutions by robertb, stanbon:
Answer by robertb(5830)
Answer by stanbon(75887)

find the slope of the graph of the linear function "f". f(-3)=-9,f(3)=9
------------
You have two points: (-3,-9) and (3,9)
-----
slope = (9--9)/(3--3) = 18/6 = 3
======================================
cheers,
Stan H.